OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

(2 square root of 2, 225°), (-2 square root of 2, 45°) (2 square root of 2, 135°), (-2 square root of 2, 315°) (2 square root of 2, 315°), (-2 square root of 2, 135°) (2 square root of 2, 45°), (-2 square root of 2, 225°)

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

Ok, we're given a rectangular coordinate this time: (2,-2), in form (X,Y) To convert to polar, use Pythagorean Theorem for r: \[r = \sqrt{X^2 + Y^2}\]

OpenStudy (anonymous):

k so r=√4+-4

OpenStudy (ipwnbunnies):

Almost. Y is -2, but when you square -2, it becomes positive 4.

OpenStudy (anonymous):

thats what i meant aha oops. So r=√8

OpenStudy (ipwnbunnies):

Yes, which simplies to 2*sqrt(2)

OpenStudy (ipwnbunnies):

Ok, for the angle. We use inverse tangent, or arctangent. \[\tan^{-1}(\frac{Y}{X}) = \theta\]

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

what do I do now?

OpenStudy (ipwnbunnies):

This is another special angle. arctan(-2/2) = arctan(-1) = 7pi/4, since the point (2,-2) is in Quadrant 4.

OpenStudy (anonymous):

oh ok ya that makes sense

OpenStudy (ipwnbunnies):

Good, we have our regular point then. (2,-2) becomes (2*sqrt(2) , 7pi/4)

OpenStudy (anonymous):

but how do we get that into the degree?

OpenStudy (ipwnbunnies):

Oh, we have to convert it to degrees. But only one of those choices has that as an answer too.

OpenStudy (ipwnbunnies):

When you convert the angle to degrees, you'll see. Btw, when converting radians to degree, multiply by 180/pi

OpenStudy (anonymous):

315 degrees, so the answer is (2√2,315 degrees), (-2√2,135 degrees)?

OpenStudy (ipwnbunnies):

Yes. Remember when we were talking about the negative r values? 135 degrees is 180 degrees less than 315. So, 135 is exactly opposite of 315. Since r is negative 2*sqrt(2), the point will be traced in the direction opposite of 135, which is 315.